# Ambient Noise Full Waveform Inversion with Neural Operators

### v3

@@ 2.3 Synthetic Tests
Before the inversion, we compute the sensitivity kernels for the HNO-AD method, following the definition in Tromp et al. (2005). In Figure 3, we put sources and receivers at nodes on the free surface and sum the gradients of the misfit with respect to the velocity parameters over all source-receiver pairs. The misfit is defined as the squared error between the HNO-predicted data with a 1D initial model and the SALVUS-simulated data with the true model (CVM-S4.26). The 1D initial model is obtained by horizontally averaging the CVM-S4.26 profile. We average the sensitivity kernels horizontally to obtain functions of depth. We can see that the sensitivity drops dramatically with depth, which indicates that updates to deeper parts of the velocity models are less reliable. Therefore, we will only present the inversion results for the top 101010 km hereafter.
@@
In the inversion stage, we freeze the HNO parameters and update $V_{P}$ and $V_{S}$ simultaneously with gradients calculated from AD. The inversion is also performed in the frequency domain, and the objective function is defined as the mean squared error (MSE) between the HNO-predicted and SALVUS-simulated data. For use with ambient noise data where the amplitude is often not well-defined, we normalize the amplitude to unity before computing the MSE. This is done by dividing the frequency-domain data, a complex-valued quantity, by its amplitude. We add no regularization term to the objective function but implicitly impose some regularization by smoothing the gradients with a Gaussian filter (Tape et al., 2010; Zhu et al., 2015). The smoothing radius is set to 3×3$\times$3 × grid spacing (945945945 m) in both horizontal and vertical directions. We average the CVM-S4.26 profile horizontally to get a 1D initial model. We start with lower frequency data and gradually feed in higher frequency components (from 0.10.10.1 to 0.50.50.5 Hz), which is also common practice in traditional FWI to mitigate cycle-skipping (Wang et al., 2018; Zhang et al., 2018). This is because skipping one cycle in lower frequency data requires larger velocity variation, whereas velocity variation is smaller at greater depth — where lower frequency data has better sensitivity. Also, the neural operator performs more accurately for lower frequencies (Zou et al., 2024). Figure 4 shows the MSE for the full frequency band in the inversion process. We use an Adam optimizer with a learning rate of 0.050.050.05 and a scheduler that decays the learning rate by half every 555 epochs. The gradients are summed over a mini-batch of 646464 data points, where a single data point corresponds to a single frequency and a single source.
@@
Figure 4: Overall MSE (mean squared error) in the FWI process, where higher frequency data is progressively incorporated. We use an Adam optimizer with a learning rate of 0.050.050.05 and a scheduler that decays the learning rate by half every 555 epochs. The batch size is 646464.
@@
We conduct FWI experiments on both noise-free data and data perturbed with different levels of Gaussian random noise. We activate 234234234 sources along the free surface of the CVM-S4.26 slice for SB1, located at the nodes. The receivers are also placed at the nodes. Figure 5 shows a sample shot of synthetic data perturbed with different levels of noise. The noise is generated from a zero-mean Gaussian distribution with a standard deviation (SD) equal to a factor times the SD of the noise-free data. The corresponding FWI results for noise-free data and noisy data are displayed in Figure 6. The noise-free inversion in the second row clearly reveals the basin model, demonstrating the forward modeling accuracy of the HNO from another perspective. The deeper part of the model is not well resolved because the source-receiver geometry and frequency content limit sensitivity to near-surface regions. When 5×5$\times$5 × noise is added, the FWI results maintain good quality, even though the seismic signal already becomes unrecognizable to the eye. This could be explained by the fact that FNOs are global operators, and the zero-mean noise might be partially averaged out in the kernel. However, real noise typically does not behave this nicely. We increase the noise level until the inversion is on the verge of collapse, with a perturbation of 10×10$\times$10 × noise that totally obliterates the waveforms.
@@
These synthetic tests serve to verify the generalizability of the neural operator trained with random fields and noise-free data to realistic velocity structures and noisy data, supervised by some definition of ground truth that is unavailable in real data. They show the excellent performance of the HNO-AD method in low signal-to-noise ratio (SNR) situations. Additionally, the entire FWI process can be completed in minutes. In contrast, FWI with the adjoint-state method takes 161616 hours for 101010 iterations and makes little progress with the same 1D initialization even for noise-free data (Figure S2). We attribute the main difference to the use of mini-batch Adam optimization in the HNO-AD method, which is not yet supported by SALVUS.
2.4 Application to Real Seismic Data from the Broader Los Angeles Region
We apply the trained neural operator to a real-world ambient noise tomography example. The real data comes from the BASIN survey that deployed 101010 linear nodal arrays in the San Gabriel (SG), Chino, and San Bernardino (SB) basins north of LA (Clayton et al., 2019). A total of 758758758 Fairfield ZLand three-component nodes with a corner frequency of 555 Hz recorded ambient noise for approximately a month. Each line consisted of 141414 to 260260260 geophones, spaced approximately 0.250.250.25 km apart. The geometry of the survey was designed to detect subsurface variations up to 0.50.50.5 km in horizontal scale. Figure 7 shows a map of the study area and station configuration. The sedimentary formations date from the opening of the LA basin in the Miocene (Wright, 1991). Details in the geologic and tectonic setting can be found in Villa et al. (2023) and Ghose et al. (2023).
Figure 7: Map of the northern Los Angeles basins, adapted from Zou & Clayton (2024). The study region is outlined by the white polygon. Triangles in different colors show nodal lines deployed in different time periods.
@@ 2.4 Application to Real Seismic Data from the Broader Los Angeles Region
where $G_{AB}(t)$ is the displacement EGF between two stations $A$ and $B$, and $C_{AB}(t)$ is the cross-correlation between the seismograms recorded at these two stations. The positive and negative lags of $C_{AB}$ are averaged to enhance the SNR and to mitigate the effect of inhomogeneous source distribution (Lin et al., 2008). Although more effective methods for addressing non-diffuse noise fields have been explored (Liu, 2020; Liu et al., 2023a), this is not the focus of this study, and we follow the standard procedure. Our cross-correlation workflow is based on Bensen et al. (2007) and Li et al. (2023b). We correlate both daytime and nighttime recordings from the one-month deployment, since using only nighttime data merely reduced the anthropogenic noise in a minor way (Li et al., 2023b). One-hour data segments are correlated and stacked to derive the final correlation. We convert the three-component data from the vertical-north-east (ZNE) to the vertical-radial-transverse (ZRT) coordinate system. The particle motion of Rayleigh waves lies in the ZR plane, whereas Love waves involve mainly horizontal motion in the transverse direction. We demean the raw seismograms and apply a bandpass filter between 0.10.10.1 and 222 Hz. Prior to the correlation, we perform temporal normalization to suppress the effects of earthquakes and apply spectral whitening to broaden the effective bandwidth.
@@
The EGF obtained from cross-correlation accounts for wave physics in a 3D medium, to which a phase shift of $\pi/4$ must be applied prior to a 2D inversion (Forbriger et al., 2014; Schäfer et al., 2014; Zhang et al., 2018). Figure S3 shows the effect of this transformation. A factor for amplitude correction related to geometrical spreading is further needed (Zhang et al., 2018), but we focus solely on the phase information, as the amplitude recovered in ANT is not well-defined. We have verified through synthetic tests that the phase information is sufficient to reconstruct the velocity models (Figure 6). We eliminate the effect of amplitude differences between 2D and 3D by normalizing the amplitude to unity during inversion. This also helps to bridge the gap between the synthetic Green’s function (SGF) and EGF, as they have different source bandwidths. We match the Z-Z and Z-R cross-correlations with the vertical and horizontal waveforms predicted by the HNO, respectively. We mainly retrieve Rayleigh waves with particle motion confined in a 2D plane. In this study, we follow the principle of Green’s function retrieval, while Tromp et al. (2010), Sager et al. (2018, 2020), and Tsai et al. (2024) established the full waveform ambient noise inversion technique that goes beyond that by interpreting correlation functions as self-consistent observables.
@@
We maintain most settings from the previous synthetic FWI results in carrying out the real data experiments, except we include an additional regularization term that encourages the inverted $V_{P}$ and $V_{S}$ to respect Brocher (2005):
@@ 3 Results
We begin the real data FWI with the longest line, SB1, which extends about 808080 km from west to east. Beneath SB1 lie two significant basins — the deeper SG basin in the west and the shallower Chino basin in the east. The depth of the sediment-basement interface has been constrained by Villa et al. (2023). In addition to this 3D basin depth model and the CVM-S4.26 (Lee et al., 2014), another reference is from Li et al. (2023b) who constructed a 3D shear wave velocity model for the same area through surface wave dispersion analysis. We start the inversion with a 1D model calculated from this reference model, which results in Figure 8a. Overall, the inverted velocity models reveal basin structures that align well with Villa et al. (2023), capturing significant features including the deeper and more heterogeneous SG basin in the west and the low-velocity zone at the edge of the relatively flat Chino basin adjacent to the SB basin. We demonstrate the robustness of this result through another two initial models, the smoothed models from Li et al. (2023b) (Figure 8b) and CVM-S4.26 (Figure 8c). Results from these different initial models show excellent overall consistency, although they differ in details. In particular, the 1D initial model can achieve performance comparable to that of models with more detailed priors.
@@
Figure 8: Inversion results for SB1 with different initial models including the (a) 1D and (b) smoothed models from Li et al. (2023b), and (c) the smoothed CVM-S4.26. The along-profile distance increases from west to east. The basin bottom from Villa et al. (2023) is delineated with black dashed lines for reference.
@@
The incorporation of higher frequency data in later stages of FWI contributes to refining the shallower structure but also introduces high-frequency noise. Some of the noise can potentially be suppressed using a signal window. Meanwhile, the inaccuracy of the neural operator in forward modeling introduces another source of uncertainty. We compare SALVUS-simulated waveforms (SGFs) using the CVM-S4.26, model from Li et al. (2023b), and our model from Figure 8a with the EGF for SB1 in Figure 9. The corresponding waveform difference is plotted in Figure S4. The point here is not to conclude that our model is better than others, but rather to demonstrate that the HNO-AD method can achieve accuracy comparable to traditional methods at a much lower cost in terms of time and effort. Also note that our model starts with a 1D initial model, while the other two models have stronger priors (Lee et al., 2014; Li et al., 2023b).
@@
Figure 9: Examples of SALVUS-simulated waveforms (SGFs) using the CVM-S4.26, model from Li et al. (2023b), and our model from Figure 8a compared with the EGF for SB1. Only $V_{S}$ is plotted in the top left for simplicity. The frequency band of the data is 0.20.20.2 to 0.50.50.5 Hz. In plotting, the waveform is normalized by trace and every third trace is displayed.
@@
The same neural operator can be directly applied to any seismic line in the study region, because it was trained to cover the longest one. While the model cannot extrapolate to larger domains than it was trained on, application to smaller domains is straightforward by simply extracting a subset of the inversion result. Additionally, no specific orientation is assumed for the 2D profiles, since the training data are generated in a Cartesian coordinate system without reference to geological orientation. This ensures that the model is not biased toward any particular direction. We show another two examples for SB4 and SG1 in Figures 10 and 11, respectively. For SB4 that crossed the Chino basin from north to south, our inversion result reveals a low-velocity zone at the southern end near the Chino fault (Figure 7), which is also seen in the CVM-S4.26 and Li et al. (2023b). The low-velocity anomalies between 0 and 151515 km present in Li et al. (2023b) are absent in both our model and CVM-S4.26. While the authors did not provide specific interpretation, the seismic line crossed the Red Hill fault at approximately 555 km, which could explain the observed discontinuity (if it is not an artifact). Interestingly, our model shows a velocity jump near the Red Hill fault. Overall, the tomography for SB4 displays a shallow (approximately 222 km) sedimentary basin in agreement with previous studies. The sediment-basement interface beneath SG1 is much deeper, as captured in all the displayed models. Our result for SG1 aligns with that for SB1 in the SG basin in terms of basin depth but captures less heterogeneity, which might be limited by the array length. These inversions can be completed in a few minutes.
@@
Figure 10: Inversion results for SB4, in comparison to the CVM-S4.26 and model from Li et al. (2023b). Only $V_{S}$ is plotted for simplicity. The along-profile distance increases from north to south. The basin bottom from Villa et al. (2023) is delineated with black dashed lines for reference.
@@
Figure 11: Inversion results for SG1, in comparison to the CVM-S4.26 and model from Li et al. (2023b). Only $V_{S}$ is plotted for simplicity. The along-profile distance increases from north to south. The basin bottom from Villa et al. (2023) is delineated with black dashed lines for reference.
@@ 4 Discussion
In this study, we have demonstrated the functionality of neural operators in realistic ambient noise full waveform inversion through a 2D case. Provided sufficient computational resources, the proposed method is expected to work in 3D without fundamental changes, while providing increased computational advantages over conventional methods. Because our model is trained within the framework of supervised learning, the fundamental assumption is that training and testing data follow the same distribution. While this assumption seldom holds strictly in practice, it serves as a reference for identifying when model predictions are likely to fail — for example, in cases where the velocity parameters deviate significantly from the training distribution. We show the training distributions of $V_{P}$ and $V_{S}$ as a function of depth in Figure 12, allowing users to decide whether to directly use our model or perform additional training. Even in cases where re-training is necessary, a pre-trained model enables transfer learning that can greatly improve learning efficiency (Yosinski et al., 2014). Another aspect is that the model is not expected to extrapolate to larger physical domains than which it was trained on (808080 km length by 404040 km depth) or to frequencies outside the preset band of interest.
@@
In performing FWI, we use gradient descent with AD instead of training a reverse model that directly maps from observed waveforms to velocity parameters. Although such end-to-end approaches are even faster (Kazei et al., 2021; Moseley et al., 2020; Wu & Lin, 2019), they provide only a single solution without supervision on the data misfit. Such approaches are arguably not ideal for inverse problems because they are ill-posed and non-unique, unless they are combined with uncertainty quantification methods. We define the misfit function for inversion as the MSE in the frequency domain for convenient use with HNO, but it would be feasible and worthwhile to explore different misfit functions in either the frequency or time domain (Bozdağ et al., 2011; Métivier et al., 2018; Fournier & Oldenburg, 2019; Sambridge et al., 2022; Sun & Alkhalifah, 2019) to better tackle the non-uniqueness and cycle-skipping problems. It is important to note, however, that the AD method requires a differentiable loss function, or some soft approximation to address discontinuity (Liu et al., 2023b).
The rapid developments in compute capabilities and machine learning research have positioned data-driven approaches as a promising new perspective for FWI. On one hand, data-driven approaches can greatly accelerate modeling and inversion. On the other hand, they rely on data provided by physics-driven approaches and, therefore, will also benefit from any advancements in those approaches. In this study, we integrate these two types of approaches by first generating training data with a physics-driven solver and then training a data-driven model as a surrogate solver. Another way is to incorporate physics into the training loss, making it possible to remove the external solver (Li et al., 2024; Raissi et al., 2019; Rasht-Behesht et al., 2022; Ren et al., 2024; Song & Alkhalifah, 2021). While training with physics loss alone is challenging, combining it with a data loss has the potential to improve generalization ability and further benefit inversion.

### v4

@@ 3 Synthetic Tests
Before the inversion, we compute the sensitivity kernels for the HNO-AD method, following the definition in tromp2005seismic. In Figure 3, we put sources and receivers at nodes on the free surface and sum the gradients of the misfit with respect to the velocity parameters over all source-receiver pairs. The misfit is defined as the squared error between the HNO-predicted data with a 1D initial model and the SALVUS-simulated data with the true model (CVM-S4.26). The 1D initial model is obtained by horizontally averaging the CVM-S4.26 profile. We average the sensitivity kernels horizontally to obtain functions of depth. We can see that the sensitivity drops dramatically with depth, which indicates that updates to deeper parts of the velocity models are less reliable. Therefore, we will only present the inversion results for the top $10$ km hereafter.
@@
In the inversion stage, we freeze the HNO parameters and update $V_{P}$ and $V_{S}$ simultaneously with gradients calculated from AD. The inversion is also performed in the frequency domain, and the objective function is defined as the mean squared error (MSE) between the HNO-predicted and SALVUS-simulated data. For use with ambient noise data where the amplitude is often not well-defined, we normalize the amplitude to unity before computing the MSE. This is done by dividing the frequency-domain data, a complex-valued quantity, by its amplitude. We add no regularization term to the objective function but implicitly impose some regularization by smoothing the gradients with a Gaussian filter (tape2010seismic; zhu2015seismic). The smoothing radius is set to $3\times$ grid spacing ($945$ m) in both horizontal and vertical directions. We average the CVM-S4.26 profile horizontally to get a 1D initial model. We start with lower frequency data and gradually feed in higher frequency components (from $0.1$ to $0.5$ Hz), which is also common practice in traditional FWI to mitigate cycle-skipping (wang2018refined; zhang2018linear). This is because skipping one cycle in lower frequency data requires larger velocity variation, whereas velocity variation is smaller at greater depth — where lower frequency data has better sensitivity. Also, the neural operator performs more accurately for lower frequencies (zou2024deep). Figure 4 shows the MSE for the full frequency band in the inversion process. We use an Adam optimizer with a learning rate of $0.05$ and a scheduler that decays the learning rate by half every $5$ epochs. For simplicity, we use a fixed number of epochs for each frequency band, balancing accuracy and efficiency, although adapting the epoch count based on convergence may be a better option. The gradients are summed over a mini-batch of $64$ data points, where a single data point corresponds to a single frequency and a single source.
@@
Figure 4: Overall MSE (mean squared error) in the FWI process, where higher frequency data is progressively incorporated. We use an Adam optimizer with a learning rate of $0.05$ and a scheduler that decays the learning rate by half every $5$ epochs. The batch size is $64$.
@@
We conduct FWI experiments on both noise-free data and data perturbed with different levels of Gaussian random noise. We activate $234$ sources along the free surface of the CVM-S4.26 slice for SB1, located at the nodes. The receivers are also placed at the nodes. Figure 5 shows a sample shot of synthetic data perturbed with different levels of noise. The noise is generated from a zero-mean Gaussian distribution with a standard deviation (SD) equal to a factor times the SD of the noise-free data. The corresponding FWI results for noise-free data and noisy data are displayed in Figure 6. The noise-free inversion in the second row clearly reveals the basin model, demonstrating the forward modeling accuracy of the HNO from another perspective. The deeper part of the model is not well resolved because the source-receiver geometry and frequency content limit sensitivity to near-surface regions. When $5\times$ noise is added, the FWI results maintain good quality, even though the seismic signal already becomes unrecognizable to the eye. This could be explained by the fact that FNOs are global operators, and the zero-mean noise might be partially averaged out in the kernel. However, real noise typically does not behave this nicely. We increase the noise level until the inversion is on the verge of collapse, with a perturbation of $10\times$ noise that totally obliterates the waveforms.
@@
These synthetic tests serve to verify the generalizability of the neural operator trained with random fields and noise-free data to realistic velocity structures and noisy data, supervised by some definition of ground truth that is unavailable in real data. They show the excellent performance of the HNO-AD method in low signal-to-noise ratio (SNR) situations. Additionally, the entire FWI process can be completed in minutes. In contrast, FWI with the adjoint-state method takes $16$ hours for $10$ iterations and makes little progress with the same 1D initialization even for noise-free data (Figure S2). We attribute the main difference to the use of mini-batch Adam optimization in the HNO-AD method, which is not yet supported by SALVUS. We provide another generalization test in Figure S3 for a velocity model obtained from a ray-theory-based method (li2023shear), which is fully independent of the training process.
4 Application to Real Seismic Data
We apply the trained neural operator to a real-world ambient noise tomography example. The real data comes from the BASIN survey that deployed $10$ linear nodal arrays in the San Gabriel (SG), Chino, and San Bernardino (SB) basins north of LA (clayton2019exposing). A total of $758$ Fairfield ZLand three-component nodes with a corner frequency of $5$ Hz recorded ambient noise for approximately a month. Each line consisted of $14$ to $260$ geophones, spaced approximately $0.25$ km apart. The geometry of the survey was designed to detect subsurface variations up to $0.5$ km in horizontal scale. Figure 7 shows a map of the study area and station configuration. The sedimentary formations date from the opening of the LA basin in the Miocene (wright1991structural). Details in the geologic and tectonic setting can be found in villa2023three and ghose2023basin.
Figure 7: Map of the northern Los Angeles basins, adapted from Zou2024. The study region is outlined by the white polygon. Triangles in different colors show nodal lines deployed in different time periods.
@@ 4 Application to Real Seismic Data
where $G_{AB}(t)$ is the displacement EGF between two stations $A$ and $B$, and $C_{AB}(t)$ is the cross-correlation between the seismograms recorded at these two stations. The positive and negative lags of $C_{AB}$ are averaged to enhance the SNR and to mitigate the effect of inhomogeneous source distribution (lin2008surface). Although more effective methods for addressing non-diffuse noise fields have been explored (liu2020finite; liu2023ambient), this is not the focus of this study, and we follow the standard procedure. Our cross-correlation workflow is based on bensen2007processing and li2023shear. We correlate both daytime and nighttime recordings from the one-month deployment, since using only nighttime data merely reduced the anthropogenic noise in a minor way (li2023shear). One-hour data segments are correlated and stacked to derive the final correlation. We convert the three-component data from the vertical-north-east (ZNE) to the vertical-radial-transverse (ZRT) coordinate system. The particle motion of Rayleigh waves lies in the ZR plane, whereas Love waves involve mainly horizontal motion in the transverse direction. We demean the raw seismograms and apply a bandpass filter between $0.1$ and $2$ Hz. Prior to the correlation, we perform temporal normalization to suppress the effects of earthquakes and apply spectral whitening to broaden the effective bandwidth.
@@
The EGF obtained from cross-correlation accounts for wave physics in a 3D medium, to which a phase shift of $\pi/4$ must be applied prior to a 2D inversion (forbriger2014line; schafer2014line; zhang2018linear). Figure S4 shows the effect of this transformation. A factor for amplitude correction related to geometrical spreading is further needed (zhang2018linear), but we focus solely on the phase information, as the amplitude recovered in ANT is not well-defined. We have verified through synthetic tests that the phase information is sufficient to reconstruct the velocity models (Figure 6). We eliminate the effect of amplitude differences between 2D and 3D by normalizing the amplitude to unity during inversion. This also helps to bridge the gap between the synthetic Green’s function (SGF) and EGF, as they have different source bandwidths. We match the Z-Z and Z-R cross-correlations with the vertical and horizontal waveforms predicted by the HNO, respectively. We mainly retrieve Rayleigh waves with particle motion confined in a 2D plane. In this study, we follow the principle of Green’s function retrieval, while tromp2010noise, sager2018towards; sager2020global, and tsai2024towards established the full waveform ambient noise inversion technique that goes beyond that by interpreting correlation functions as self-consistent observables.
@@
We maintain most settings from the previous synthetic FWI results in carrying out the real data experiments, except we include an additional regularization term that encourages the inverted $V_{P}$ and $V_{S}$ to respect brocher2005empirical:
@@
We begin the real data FWI with the longest line, SB1, which extends about $80$ km from west to east. Beneath SB1 lie two significant basins — the deeper SG basin in the west and the shallower Chino basin in the east. The depth of the sediment-basement interface has been constrained by villa2023three. In addition to this 3D basin depth model and the CVM-S4.26 (lee2014full), another reference is from li2023shear who constructed a 3D shear wave velocity model for the same area through surface wave dispersion analysis. We start the inversion with a 1D model calculated from this reference model, which results in Figure 8a. Overall, the inverted velocity models reveal basin structures that align well with villa2023three, capturing significant features including the deeper and more heterogeneous SG basin in the west and the low-velocity zone at the edge of the relatively flat Chino basin adjacent to the SB basin. We demonstrate the robustness of this result through another two initial models, the smoothed models from li2023shear (Figure 8b) and CVM-S4.26 (Figure 8c). Results from these different initial models show excellent overall consistency, although they differ in details. In particular, the 1D initial model can achieve performance comparable to that of models with more detailed priors. The corresponding error curves are provided in Figure S5.
@@
Figure 8: Inversion results for SB1 with different initial models including the (a) 1D and (b) smoothed models from li2023shear, and (c) the smoothed CVM-S4.26. The along-profile distance increases from west to east. The basin bottom from villa2023three is delineated with black dashed lines for reference.
@@
The incorporation of higher frequency data in later stages of FWI contributes to refining the shallower structure but also introduces high-frequency noise. Some of the noise can potentially be suppressed using a signal window. Meanwhile, the inaccuracy of the neural operator in forward modeling introduces another source of uncertainty. We compare SALVUS-simulated waveforms (SGFs) using the CVM-S4.26, model from li2023shear, and our model from Figure 8a with the EGF for SB1 in Figure 9. The corresponding waveform difference is plotted in Figure S6. The point here is not to conclude that our model is better than others, but rather to demonstrate that the HNO-AD method can achieve accuracy comparable to traditional methods at a much lower cost in terms of time and effort. Also note that our model starts with a 1D initial model, while the other two models have stronger priors (lee2014full; li2023shear).
@@
Figure 9: Examples of SALVUS-simulated waveforms (SGFs) using the CVM-S4.26, model from li2023shear, and our model from Figure 8a compared with the EGF for SB1. Only $V_{S}$ is plotted in the top left for simplicity. The frequency band of the data is $0.2$ to $0.5$ Hz. In plotting, the waveform is normalized by trace and every third trace is displayed.
@@
The same neural operator can be directly applied to any seismic line in the study region, because it was trained to cover the longest one. While the model cannot extrapolate to larger domains than it was trained on, application to smaller domains is straightforward by simply extracting a subset of the inversion result. Additionally, no specific orientation is assumed for the 2D profiles, since the training data are generated in a Cartesian coordinate system without reference to geological orientation. This ensures that the model is not biased toward any particular direction. We show another two examples for SB4 and SG1 in Figures 10 and 11, respectively. For SB4 that crossed the Chino basin from north to south, our inversion result reveals a low-velocity zone at the southern end near the Chino fault (Figure 7), which is also seen in the CVM-S4.26 and li2023shear. The low-velocity anomalies between 0 and $15$ km present in li2023shear are absent in both our model and CVM-S4.26. While the authors did not provide specific interpretation, the seismic line crossed the Red Hill fault at approximately $5$ km, which could explain the observed discontinuity (if it is not an artifact). Interestingly, our model shows a velocity jump near the Red Hill fault. Overall, the tomography for SB4 displays a shallow (approximately $2$ km) sedimentary basin in agreement with previous studies. The sediment-basement interface beneath SG1 is much deeper, as captured in all the displayed models. Our result for SG1 aligns with that for SB1 in the SG basin in terms of basin depth but captures less heterogeneity, which might be limited by the array length. These inversions can be completed in a few minutes.
@@
Figure 10: Inversion results for SB4, in comparison to the CVM-S4.26 and model from li2023shear. Only $V_{S}$ is plotted for simplicity. The along-profile distance increases from north to south. The basin bottom from villa2023three is delineated with black dashed lines for reference.
@@
Figure 11: Inversion results for SG1, in comparison to the CVM-S4.26 and model from li2023shear. Only $V_{S}$ is plotted for simplicity. The along-profile distance increases from north to south. The basin bottom from villa2023three is delineated with black dashed lines for reference.
@@ 5 Discussion
In this study, we have demonstrated the functionality of neural operators in realistic ambient noise full waveform inversion through a 2D case. Provided sufficient computational resources, the proposed method is expected to work in 3D without fundamental changes, while providing increased computational advantages over conventional methods. Because our model is trained within the framework of supervised learning, the fundamental assumption is that training and testing data follow the same distribution. While this assumption seldom holds strictly in practice, it serves as a reference for identifying when model predictions are likely to fail — for example, in cases where the velocity parameters deviate significantly from the training distribution. We show the training distributions of $V_{P}$ and $V_{S}$ as a function of depth in Figure 12, allowing users to decide whether to directly use our model or perform additional training. Even in cases where re-training is necessary, a pre-trained model enables transfer learning that can greatly improve learning efficiency (yosinski2014transferable). Another aspect is that the model is not expected to extrapolate to larger physical domains than which it was trained on ($80$ km length by $40$ km depth) or to frequencies outside the preset band of interest.
@@
In performing FWI, we use gradient descent with AD instead of training a reverse model that directly maps from observed waveforms to velocity parameters. Although such end-to-end approaches are even faster (kazei2021mapping; moseley2020deep; wu2019inversionnet), they provide only a single solution without supervision on the data misfit. Such approaches are arguably not ideal for inverse problems because they are ill-posed and non-unique, unless they are combined with uncertainty quantification methods. We define the misfit function for inversion as the MSE in the frequency domain for convenient use with HNO, but it would be feasible and worthwhile to explore different misfit functions in either the frequency or time domain (bozdaug2011misfit; metivier2018optimal; fournier2019inversion; sambridge2022geophysical; sun2019application) to better tackle the non-uniqueness and cycle-skipping problems. It is important to note, however, that the AD method requires a differentiable loss function, or some soft approximation to address discontinuity (liu2023robust).
The rapid developments in compute capabilities and machine learning research have positioned data-driven approaches as a promising new perspective for FWI. On one hand, data-driven approaches can greatly accelerate modeling and inversion. On the other hand, they rely on data provided by physics-driven approaches and, therefore, will also benefit from any advancements in those approaches. In this study, we integrate these two types of approaches by first generating training data with a physics-driven solver and then training a data-driven model as a surrogate solver. Another way is to incorporate physics into the training loss, making it possible to remove the external solver (li2024physics; raissi2019physics; Rasht‐Behesht; ren2024seismicnet; SongAlkhalifah). While training with physics loss alone is challenging, combining it with a data loss has the potential to improve generalization ability and further benefit inversion.